# Regularized Q-Learning with Linear Function Approximation

### v1

@@ Abstract
Several successful reinforcement learning algorithms make use of regularization to promote multi-modal policies that exhibit enhanced exploration and robustness. With functional approximation, the convergence properties of some of these algorithms (e.g. soft Q-learning) are not well understood. In this paper, we consider a single-loop algorithm for minimizing the projected Bellman error with finite time convergence guarantees in the case of linear function approximation. The algorithm operates on two scales: a slower scale for updating the target network of the state-action values, and a faster scale for approximating the Bellman backups in the subspace of the span of basis vectors. We show that, under certain assumptions, the proposed algorithm converges to a stationary point in the presence of Markovian noise. In addition, we provide a performance guarantee for the policies derived from the proposed algorithm.
{IEEEkeywords}
Linear function approximation, Q-learning, reinforcement learning.
@@ 1 Introduction
[…]
Numerous reinforcement learning algorithms have achieved remarkable success by employing some form of regularization [1, 2]. The specific choice of regularization choices induces a specific trade-off between exploration and exploitation. For example, negative Shannon entropy regularization has been successfully employed in [3, 2]. In that case, the learned policies are robust to perturbations in both the dynamics and the reward function [4]. Regularization can also be used to impose sparsity of learned policies (e.g., Tsallis entropy [5, 6, 7]). This approach maintains a multi-modal policy form but restricts the support to only optimal and near-optimal actions, thereby mitigating the impact of detrimental actions. Recent studies [8, 7] develop the framework for analyzing regularized Markov Decision Processes (MDPs) with a broad class of regularizers. The regularized MDPs are proven to exhibit robustness in scenarios where the reward function is uncertain [9], and they boast a faster convergence rate during policy optimization [10, 11, 12, 13]. In this study, we investigate the convergence properties of an off-policy algorithm tailored for approximating the optimal value function in MDPs equipped with strongly convex regularizers, within the context of linear function approximation.
Q-learning, as presented in [14], is a model-free, off-policy reinforcement learning algorithm. While it has demonstrated convergence in tabular settings [15], its convergence performance degrades with larger discount factors. Recently, a novel variant termed ‘relative Q-learning’ was introduced in [16]; it is designed to provide uniform bounds on sample complexity, regardless of the discount factor. However, the convergence of Q-learning with linear function approximation has not been formally established except under restrictive assumptions, such as an exceedingly small discount factor [17, 18]. Generally, the function obtained by applying the Bellman operator to a function within the span of predefined features cannot be expressed as a linear combination of these features [19]. The Bellman operator exhibits contraction properties with respect to the $\ell_{\infty}$-norm, and the projection onto the span of predefined features is non-expansive with respect to the $\ell_{2}$-norm. However, when these operators are composed to form the projected Bellman operator, the resulting operator does not necessarily retain the contraction property with respect to any norm. As a consequence, divergence may occur during the iteration process [20]. A substantial body of work has been devoted to developing variants of the Q-learning algorithm to overcome this issue. [21, 22] demonstrate convergence by employing ridge regularization on the approximator (also referred to as the main network); however, these approaches solve an MDP with a decreased discount factor instead of the original problem. [23] demonstrated asymptotic convergence, and further extended the framework to a linear function approximation scheme subject to specific conditions on the features. Introducing a target network, which is a delayed copy of the main network, has been shown to effectively stabilize the off-policy learning process in deep reinforcement learning [24]. Within Coupled Q-learning [25], the target network undergoes updates at a slower rate instead of copying the main network. Although the convergence of the aforementioned methods is established, the quality of the convergent points is ensured only with a sufficiently low discount factor. Zap Zero algorithm [26] approximates the Newton-Raphon flow without matrix inversion and is guaranteed to converge to the optimal solution under mild assumptions. Greedy Gradient Q-learning (Greedy-GQ) with linear function approximation is a provably convergent algorithm under mild assumptions [27]. The algorithm utilizes gradient descent and seeks to minimize the Mean-Square Projected Bellman Error (MSPBE), which is non-convex in general. The finite-time convergence guarantees to the stationary points were established recently in [28, 29, 30, 31]. Recent work [32] demonstrates that employing a target network, in conjunction with the truncation operator on Bellman backup values, effectively confines the estimated value function within a desirable neighbor of the optimal value function.
1.1 Our Contributions
We have developed an algorithm with finite-time guarantees for learning the solution to a broad class of regularized Markov Decision Processes (MDPs) using linear function approximation. Our contributions are as follows:
  • We introduce a smooth truncation operator that is applied to the Bellman backup in regularized MDPs, preserving the smoothness of the regularized Bellman operator.
  • We propose a bi-level formulation for regularized Q-learning with target (upper level) and main (lower level) solution (or network), respectively. In this approach, the lower-level problem is to find the projected Bellman backup for the target network – this is a strongly convex problem. The upper-level problem aims to determine the optimal target network parameter $\theta^{*}$ that minimizes the non-convex Mean-Square Projected Bellman Error (MSPBE), denoted as $J(\theta)$, which incorporates a smooth truncation operator. This formulation serves as a foundation for the development of the proposed algorithm.
  • We develop a single-loop algorithm to learn the solution of the proposed bi-level formulation in the presence of Markovian noise. Notably, we demonstrate that the algorithm converges to a stationary point at a rate $T^{-1}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla J(\theta^{t})\|=\mathcal{O}(T^{-1/4})$. Moreover, we establish finite-time performance guarantees for the learned policies. These guarantees reveal that, compared to the optimal policy, the expected performance gap of policies obtained in $T>0$ iterations is $\mathcal{O}(T^{-1/4})$, augmented by the function approximation error and the error introduced by the smooth truncation operator.
  • We carry out numerical experiments to illustrate the impact of errors introduced by the smooth truncation operator on the solution quality. Additionally, our evaluation of policies derived from the proposed algorithms reveals superior performance compared to existing methods in most scenarios.
@@ 1.2 Related Work
A large number of studies have been done on variants of the Q-learning algorithm within the framework of function approximation. In this subsection, we briefly discuss some works that are most closely related to our work.
@@
The stability of Q-learning and its related variants, particularly in the context of function approximation, has long been a focus of research [33, 34]. Numerous techniques have been proposed to address these stability concerns, including the target network [24], double estimator [35], fitted value iteration [36], gradient-based approaches [27], and stochastic Newton-Raphson [37], among others. Furthermore, ensuring the convergence of these variant algorithms remains a more significant challenge in the field of reinforcement learning, even when employing linear function approximation [32, 33].
Zap Q-learning [38, 37] and its variant, Zap Zero [26], mimic the classical Newton-Raphson algorithm. These algorithms guarantee asymptotic convergence under mild assumptions, though they come with substantial computational costs due to matrix inversion or its approximation process.
@@
Greedy-GQ [27], inspired by the success of TDC in off-policy evaluation [19], employs a two-timescale gradient-based approach to minimize the MSPBE, and its asymptotic convergence to a stationary point is established over i.i.d. sampled data. Recent work [30] provides a finite-time guarantee for the convergence of Greedy-GQ under Markovian noise with a rate of $\mathcal{O}(T^{-1/3}\log T)$, and it has been improved to $\mathcal{O}(T^{-1/2}\log T)$ in [29], albeit without providing guarantees for the boundedness of the iterates. Subsequent work [31] further improves the result to $\mathcal{O}(T^{-1/2})$ with the use of a mini-batch gradient. It should be noted that the criteria they employ, in terms of rate, diverge from ours, given their consideration of the square of the gradient’s norm. While our approach bears similarities to Greedy-GQ in its formulation, the objectives differ. Greedy-GQ’s lower-level problem seeks to approximate the TD error, whereas our method focuses on maintaining a main network to approximate Bellman backups within the span of the basis. Moreover, we utilize the truncation and projection operators to ensure the boundedness of the iterates without compromising the convergence rate.
Employing ridge regularization in Q-learning allows for convergence guarantees with linear function approximation, without imposing stringent assumptions [21, 22]. However, this approach diverges from addressing the original problem, as it inherently reduces the discount factor. In contrast, our work views the regularized MDP as the original problem in which it is the policy, not the value function, that is regularized. Our proposed updates the main network in the same manner as in Coupled Q-learning [25], while the target network in Coupled Q-learning is updated toward the fixed point of the composed operator constructed by combining the ”un-normalized” orthogonal projection and the Bellman operator. Consequently, unlike our algorithm which focuses on minimizing a specific objective function, Coupled Q-Learning does not target the minimization of any particular objective. While Coupled Q-learning converges asymptotically without stringent assumptions, its estimated value function proves meaningful only with a low discount factor.
Over-parameterized neural networks serve as non-linear function approximators and have found extensive applications in various reinforcement learning algorithms. These include TD-learning [39], the policy gradient method [40], and Q-learning [41, 42]. Furthermore, recent studies have established finite-sample bounds for these algorithms [43, 44, 45]. However, over-parameterization typically necessitates a neural network with a sufficiently large width, a scenario in which the functions in the considered function class are nearly linear around the initialization point [46]. In contrast, the linear function approximation employed in this study generally operates under the ‘under-parameterized’ framework, which can not only lower computational costs in practice but also yield deeper insights into under-parameterized cases. More importantly, neural network-based estimations are less interpretable compared to those using linear function approximation. In the latter case, feature selection is usually driven by an understanding of the model and specific interests.
@@ 2.1 Regularized Markov Decision Process
We consider an infinite-horizon regularized Markov Decision Processes (MDPs) with finite action space. An MDP can be described as a tuple $\mathcal{M}\triangleq\left(\mathcal{S},\mathcal{A},P,\mu_{0},R,\gamma\right)$, where $\mathcal{S}$ denotes the state space and $\mathcal{A}$ denotes the action space. $P$ is the dynamics: $P(s^{\prime}\mid s,a)$ is the transition probability from state $s$ to $s^{\prime}$ by taking action $a$. $\mu_{0}$ is the initial distribution of the state. The reward function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, is bounded in absolute value by $R_{\max}\geq 0$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. $\gamma\in(0,1)$ is the discount factor. A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is defined as a mapping from the state space $\mathcal{S}$ to the probability distribution over the action space $\mathcal{A}$. For discrete time $t\geq 0$, the trajectory, starting from an initial state $s_{0}\sim\mu_{0}(\cdot)$, generated by the policy $\pi$ in the MDP $\mathcal{M}$ can be represented as a set of transition tuples $\left\{s_{t},a_{t},s_{t}^{\prime}\right\}_{t\geq 0}$, where $a_{t}\sim\pi(\cdot\mid s_{t})$, $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$, and $s_{t+1}=s_{t}^{\prime},\forall t\geq 0$.
@@
Let $G:\Delta(\mathcal{A})\rightarrow\mathbb{R}$ be a strongly convex bounded function, where there exists a constant $B>0$ such that $|G(p)|\leq B$ for all $p\in\Delta(\mathcal{A})$. Then the convex conjugate of $G^{*}:\mathbb{R}^{|\mathcal{A}|}\rightarrow\mathbb{R}$ is given by
$G^{*}(q)=\max_{p\in\Delta(\mathcal{A})}\left\langle p,q\right\rangle-G(p),\quad\forall q\in\mathbb{R}^{|\mathcal{A}|}.$
One commonly used choice of $G$ is negative Shannon entropy; in that case $G(p)=\left\langle p,\log p\right\rangle$, and the convex conjugate is $G^{*}(q)=\log\left(\left\langle\exp(q),\boldsymbol{1}\right\rangle\right)$, where $\boldsymbol{1}$ is the vector of ones. Another example is a specific form of negative Tsallis entropy, $G(p)=\frac{1}{2}(\left\|p\right\|^{2}-1)$, which leads to a sparse MDP [5]. For the sake of clarity, we define $G_{\tau}\triangleq\tau\cdot G$, where $\tau>0$ is a positive coefficient that determines the degree of regularization applied. Consequently, the convex conjugate of $G_{\tau}$ is
$G_{\tau}^{*}(q)=\max_{p\in\Delta(\mathcal{A})}\left\langle p,q\right\rangle-G_{\tau}(p)=\tau\cdot G^{*}(q/\tau),\quad\forall q\in\mathbb{R}^{|\mathcal{A}|}.$
Unlike the unregularized MDP, the objective of a regularized MDP $\mathcal{M}_{\tau}$ is to maximize not only the expected cumulative reward but also include a policy regularization term [7]. The regularized state-action value function under a given policy $\pi$, starting from a specific state-action pair $(s,a)$, is defined as
$Q_{\pi,\tau}(s,a)=R(s,a)\\ +\mathbb{E}\bigg{[}\sum_{t=1}^{\infty}\gamma^{t}\big{(}R\left(s_{t},a_{t}\right)-G_{\tau}(\pi(\cdot\mid s_{t}))\big{)}\bigg{|}s_{0}=s,a_{0}=a,P,\pi\bigg{]}.$
In a similar manner, the regularized state value function can be defined as
$V_{\pi,\tau}(s)=$ $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(R\left(s_{t},a_{t}\right)-G_{\tau}(\pi(\cdot\mid s_{t}))\right)\bigg{|}s_{0}=s,P,\pi\right]$
$=$ $\mathbb{E}_{a\sim\pi(\cdot\mid s)}\left[Q_{\pi,\tau}(s,a)\right]-G_{\tau}\left(\pi\left(\cdot\mid s\right)\right);$ (1)
this quantity represents the expected cumulative sum of rewards and regularization terms, commencing from state $s$ and adhering to policy $\pi$ subsequently, within the MDP $\mathcal{M}_{\tau}$ governed by dynamics $P$.
Some properties of regularized MDPs are stated in the following proposition. For a detailed analysis, see [8].
Proposition 2.1
@@
Let $G$ be a strongly convex function bounded by $B>0$ and $\tau>0$ be a coefficient associated with $G$. The following hold:
  1. 1.
    The gradient of the convex conjugate with respect to the input $q$ is $\nabla G_{\tau}^{*}(q)=\arg\max_{p\in\Delta(\mathcal{A})}\left\langle p,q\right\rangle-G_{\tau}(p),\forall q\in\mathbb{R}^{|\mathcal{A}|}$. $\nabla G_{\tau}^{*}$ is $\frac{L_{G}}{\tau}$-Lipschitz, i.e., there exists a constant $L_{G}>0$ such that $\left\|\nabla G_{\tau}^{*}(q_{1})-\nabla G_{\tau}^{*}(q_{2})\right\|\leq\frac{L_{G}}{\tau}\left\|q_{1}-q_{2}\right\|,\forall q_{1},q_{2}\in\mathbb{R}^{|\mathcal{A}|}$.
  2. 2.
    The conjugate function is bounded: $|G^{*}_{\tau}\left(q\right)-\max_{i}q_{(i)}|\leq\tau B,\forall q\in\mathbb{R}^{|\mathcal{A}|}$, where $q_{(i)}$ is the $i$-th element of the vector $q$.
  3. 3.
    For any policy $\pi$, the regularized value functions $V_{\pi,\tau}$ and $Q_{\pi,\tau}$ satisfy regularized Bellman equation, for all $(s,a)\in\mathcal{S}\times\mathcal{A}$:
    $Q_{\pi,\tau}(s,a)=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P\left(\cdot\mid s,a\right)}\left[V_{\pi,\tau}\left(s^{\prime}\right)\right].$
  4. 4.
    The optimal state value function $V^{*}_{\tau}(s)=\max_{\pi}V_{\pi,\tau}(s)=G^{*}_{\tau}\left(Q^{*}_{\tau}\left(s,\cdot\right)\right),\forall s\in\mathcal{S}$, where $Q^{*}_{\tau}(\cdot,\cdot)$ is the optimal state-action value function and, for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, it satisfies
    $Q^{*}_{\tau}(s,a)=$ $R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P\left(\cdot\mid s,a\right)}\left[V^{*}_{\tau}\left(s^{\prime}\right)\right].$
    This equation is commonly referred to as the optimal regularized Bellman equation.
  5. 5.
    The optimal policy $\pi^{*}_{\tau}(\cdot\mid s)=\nabla G^{*}_{\tau}\left(Q^{*}_{\tau}\left(s,\cdot\right)\right)$, for all $s\in\mathcal{S}$.
Given properties in Proposition 2.1, we define the regularized Bellman operator $\mathcal{B}_{\tau}$ as, for an arbitrary mapping $X:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$:
$\mathcal{B}_{\tau}X(s,a)\triangleq R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\left[G^{*}_{\tau}\left(X\left(s^{\prime},\cdot\right)\right)\right],$
and it is $\gamma$-contraction in $\ell_{\infty}$-norm [8]. Observe that the optimal state-action value function $Q^{*}_{\tau}$ is the unique fixed point of the optimal regularized Bellman operator: $Q^{*}_{\tau}=\mathcal{B}_{\tau}Q^{*}_{\tau}$. This fixed point can be derived through value iteration.

### v3

@@ Abstract
Regularized Markov Decision Processes serve as models of sequential decision making under uncertainty wherein the decision maker has limited information processing capacity and/or aversion to model ambiguity. With functional approximation, the convergence properties of learning algorithms for regularized MDPs (e.g. soft Q-learning) are not well understood because the composition of the regularized Bellman operator and a projection onto the span of basis vectors is not a contraction with respect to any norm. In this paper, we consider a bi-level optimization formulation of regularized Q-learning with linear functional approximation. The lower level optimization problem aims to identify a value function approximation that satisfies Bellman’s recursive optimality condition and the upper level aims to find the projection onto the span of basis vectors. This formulation motivates a single-loop algorithm with finite time convergence guarantees. The algorithm operates on two time-scales: updates to the projection of state-action values are ‘slow’ in that they are implemented with a step size that is smaller than the one used for ‘faster’ updates of approximate solutions to Bellman’s recursive optimality equation. We show that, under certain assumptions, the proposed algorithm converges to a stationary point in the presence of Markovian noise. In addition, we provide a performance guarantee for the policies derived from the proposed algorithm.
@@ 1 Introduction
Recent literature has examined regularized Markov Decision Processes MDPs as models of sequential decision making under uncertainty that account for limited information processing capacity [1, 2, 3] and aversion to model misspecification or ambiguity [4]. In this literature, low entropy policies are not necessarily optimal because the agent has limited capacity to process new information and/or is averse to model uncertainty. A parallel strand of the literature in reinforcement learning algorithms [5, 6, 7, 8] has also employed equivalent formulations with different motivations, namely that of ensuring learned policies are robust to perturbations in both the dynamics and the reward function [9]. For example, negative Shannon entropy regularization has been employed in [10, 6] to motivate “soft” Q-learning and actor-critic algorithms with robust performance. Regularization has also been used to ensure learned policies are sparse (e.g., Tsallis entropy [11, 12, 8]). This approach maintains a multi-modal policy form but restricts the support to only optimal and near-optimal actions, thereby mitigating the impact of detrimental actions. Nonetheless, the convergence of regularized Q-learning with linear functional approximation is not guaranteed in general [13]. This is because the composition of a projection onto the span of predefined features with the regularized Bellman operator is not necessarily a contraction with respect to any norm. Hence, a regularized Q-learning iteration process may diverge because the objectives of maximizing expected discounted reward and projecting onto the linear basis may “push and pull” solution updates in significantly different directions. To our knowledge, there are no algorithms with finite-time guarantees to solve regularized MDP models with linear function approximation and general strongly convex and bounded regularizer.
To address this gap, we consider a bi-level optimization approach for regularized $Q$-learning that enables us to overcome this challenge. The lower level optimization problem aims to identify a value function approximation (referred to as the main solution) that satisfies Bellman’s recursive optimality condition. The upper level aims to find a projection onto the span of basis vectors, i.e. a parametrized linear function (referred to as the target solution) that best approximates the main solution in the sense of minimizing mean square error. This formulation motivates a single-loop algorithm wherein at each iteration both the upper level and the lower level solutions are updated with different stepsizes: updates to the projection of state-action values are ‘slow’ in that they are implemented with a step size that is smaller than the one used for ‘faster’ updates of approximate solutions to Bellman’s recursive optimality equation. We demonstrate that in the presence of Markovian noise the proposed algorithm converges to a stationary point at a rate $\mathcal{O}(T^{-1/4}(\log T)^{1/2})$ with $T>0$ denoting the total number of iterations. This result is cast in terms of the $\ell_{2}$ norm rather than the squared $\ell_{2}$ norm, because unlike other single-loop algorithms for reinforcement learning [14] the proposed algorithm does not require a projection for implementing upper level updates. Lastly, we establish finite-time performance guarantees for the learned policies. These guarantees reveal that, compared to the optimal policy, the expected performance gap of policies is $\mathcal{O}(T^{-1/4}(\log T)^{1/2})$, augmented by the function approximation error and the error introduced by the smooth truncation operator.
@@ 1.1 Related Work
The stability of Q-learning and its related variants, particularly in the context of function approximation, has long been a focus of research [15, 16]. Numerous techniques have been proposed to address these stability concerns, including the target network [17], double estimator [18], fitted value iteration [19], gradient-based approaches [20], Zap Q-learning [21, 22] and its variant, Zap Zero [23] among others. In this subsection, we briefly discuss some works that are most closely related to our work.
The single-loop algorithm proposed in this paper stands in contrast with nested-loop algorithms, e.g. [24, 25] which require greater computational effort unless precise fine-tuning of the number of inner-loop iterations is available. Coupled Q-learning [26] is a single-loop approach that also relies on updates to the main and target solutions with decreasing step-sizes in a two-time scale manner as in [27]. However, as shown in [24] (see appendix E.3) the performance guarantee for Coupled Q-learning involves significant bias unless the discount factor is sufficiently small.
Our work is also related to approaches to Q-learning with ridge regularization [28, 29]. However, ridge regularization results in a modified Bellman equation with an implicitly scaled down discount factor (see appendix E.2 in [24]). Therefore it is not clear how the identified solution approximates the solution of either the original MDP or its regularized version. In contrast, the finite-time guarantee provided in this paper ensures the identified solution approximately minimizes Mean Squared Projected Bellman Error (MSPBE), a measure of the extent to which the identified solution fails to satisfy Bellman’s optimality equation.
@@
Finally, the Greedy-GQ algorithm [20], inspired by the success of TDC in off-policy evaluation [30], is arguably the closest to our work. Greedy-GQ employs a two-timescale gradient-based approach to minimize the MSPBE, and its asymptotic convergence to a stationary point is established over i.i.d. sampled data. Recent work [31] provides a finite-time guarantee for the convergence of Greedy-GQ under Markovian noise with a rate of $\mathcal{O}(T^{-1/3}\log T)$, and it has been improved to $\mathcal{O}(T^{-1/2}\log T)$ in [32]. Subsequent work [33] further improves the result to $\mathcal{O}(T^{-1/2})$ with the use of a mini-batch gradient.
The performance guarantees studied in these Greedy-GQ papers concern the solution to a standard MDP problem (i.e. not regularized). In contrast, in this paper, we consider the solution of regularized MDP problem wherein the regularizer belongs to a general class of strongly convex and bounded functions. This class includes negative Shannon and negative Tsallis entropy functions. Such regularized MDPs are used to model sequential decision making under uncertanity with limited information processing capacity [1, 2, 3] and/or aversion to model misspecification or ambiguity [4]. To our knowledge, there are no algorithms with finite-time guarantees to solve such regularized MDP models with linear function approximation. Our work contributes to fill this gap.
@@ 2 Preliminaries
2.1 Markov Decision Processes (MDPs)
We consider an infinite-horizon regularized Markov Decision Processes (MDPs) with finite action space. An MDP can be described as a tuple $\mathcal{M}\triangleq\left(\mathcal{S},\mathcal{A},P,\mu_{0},R,\gamma\right)$, where $\mathcal{S}$ denotes the state space and $\mathcal{A}$ denotes the action space. $P$ is the dynamics: $P(s^{\prime}\mid s,a)$ is the transition probability from state $s$ to $s^{\prime}$ by taking action $a$. $\mu_{0}$ is the initial distribution of the state. The reward function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, is bounded in absolute value by $R_{\max}\geq 0$, i.e., $-R_{\max}\leq R(s,a)\leq R_{\max}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. $\gamma\in(0,1)$ is the discount factor. A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is defined as a mapping from the state space $\mathcal{S}$ to the probability distribution over the action space $\mathcal{A}$. For discrete time $t\geq 0$, the trajectory, starting from an initial state $s_{0}\sim\mu_{0}(\cdot)$, generated by the policy $\pi$ in the MDP $\mathcal{M}$ can be represented as a set of transition tuples $\left\{s_{t},a_{t},s_{t}^{\prime}\right\}_{t\geq 0}$, where $a_{t}\sim\pi(\cdot\mid s_{t})$, $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$, and $s_{t+1}=s_{t}^{\prime},\forall t\geq 0$.
The state value and state-action value functions corresponding to a policy $\pi$ are defined by the expected cumulative rewards obtained when initiating from state $s\in\mathcal{S}$ or state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, and subsequently adhering to policy $\pi$, respectively:
$V_{\pi}(s)\triangleq\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}R\left(s_{t},a_{t}\right)\bigg{|}s_{0}=s,P,\pi\big{]},$
and $Q_{\pi}(s,a)=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V_{\pi}(s^{\prime})]$. In general, the objective of an MDP is to identify the optimal policy $\pi^{*}$ that maximizes the expected cumulative discounted reward. Mathematically, this can be expressed as:
$\pi^{*}\in\arg\max_{\pi}\mathbb{E}_{s\sim\mu_{0}}\left[V_{\pi}(s)\right].$
Let $Q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. The Bellman operator is defined as:
$\mathcal{B}Q(s,a)\triangleq R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(s,a)}\big{[}\max_{a\in\mathcal{A}}\{Q(s^{\prime},a)\}\big{]}.$ (1)
Under certain assumptions, this operator is a contraction with modulus $\gamma<1$ and its unique fixed point $Q^{*}$ characterizes the optimal value function as follows:
$V_{\pi^{*}}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a),\quad\forall s\in\mathcal{S}.$
@@ 2.2 Regularized Markov Decision Process
Let $G:\Delta(\mathcal{A})\rightarrow\mathbb{R}$ be a strongly convex function, that is bounded, i.e. there exists $B>0$ such that $-B\leq G(p)\leq B$ for all $p\in\Delta(\mathcal{A})$. For a given policy $\pi$ , the regularized value function is defined as
$V_{\pi,G_{\tau}}(s)$ $\triangleq\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}\big{(}R\left(s_{t},a_{t}\right)-G_{\tau}(\pi(\cdot\mid s_{t}))\big{)}\big{|}s_{0}=s,P,\pi\big{]}$
where $G_{\tau}\triangleq\tau\cdot G$, and $\tau>0$ is a positive coefficient that determines the degree of regularization applied. Let us also define $Q_{\pi,G_{\tau}}(s,a)\triangleq R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V_{\pi,G_{\tau}}(s^{\prime})]$.
The regularized Bellman operator is defined as follows:
$\mathcal{B}_{G_{\tau}}Q(s,a)\triangleq R(s,a)\\ +\gamma\mathbb{E}_{s^{\prime}\sim P(s,a)}\big{[}\max_{\pi}\big{\{}\mathbb{E}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[Q(s^{\prime},a^{\prime})]-G_{\tau}(\pi(\cdot|s^{\prime}))\big{\}}\big{]}.$
Note that without regularization (i.e. $\tau=0$), definitions (1) and (2.2) are equivalent. The regularized Bellman operator can be re-written as:
$\mathcal{B}_{G_{\tau}}Q(s,a)=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(s,a)}\big{[}G^{*}_{\tau}(Q(s^{\prime},\cdot))\big{]}$ (2)
wherein $G^{*}_{\tau}$ is the convex conjugate of $G_{\tau}$, i.e.:
$G^{*}_{\tau}(Q(s^{\prime},\cdot))\triangleq\max_{\pi}\big{\{}\mathbb{E}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[Q(s^{\prime},a^{\prime})]-G_{\tau}(\pi(\cdot|s^{\prime}))\big{\}}$
The regularized Bellman operator is a contraction in $\ell_{\infty}$ norm with modulus $\gamma<1$ (see Proposition 2 in [7]). Let $V^{*}_{G_{\tau}}(s)\triangleq\max_{\pi}\mathbb{E}_{s\sim\mu_{0}}\big{[}V_{\pi,G_{\tau}}(s)\big{]}$. According to Theorem 1 in [7], the unique fixed point of the regularized Bellman operator, say $Q^{*}_{G_{\tau}},$ characterizes the optimal value in the following sense:
$G^{*}_{\tau}(Q^{*}_{G_{\tau}}(s,\cdot))=\max_{\pi}V_{\pi,G_{\tau}}(s)$
Some Examples of Regularizers: A common choice for regularizer is negative Shannon entropy. As mentioned before, this choice has been used as a way to account for information processing costs [1, 2, 3] and to motivate the “soft” Q-learning algorithm. In this case $G_{\tau}(\pi(\cdot|s^{\prime}))=\tau\sum_{a\in A}\pi(a|s^{\prime})\log\pi(a|s^{\prime})$ and
$G^{*}_{\tau}(Q(s^{\prime},\cdot))=\tau\log\big{(}\sum_{a\in A}\exp{\frac{1}{\tau}Q(s^{\prime},a)}\big{)}$
Another choice of regularizer is the negative Tsallis entropy, $G(\pi(\cdot|s^{\prime}))=\frac{1}{2}(\left\|\pi(\cdot|s^{\prime})\right\|^{2}-1)$. In [11], the authors show that this regularizer induces a sparse and multi-modal optimal policy. Some additional properties of regularized MDPs are stated in the following proposition (see Proposition 1 in [7] for proofs).
Proposition 2.1.
@@
Let $G$ be a strongly convex function bounded by $B>0$ and $\tau>0$ be the regularization coefficient associated with $G$. The following hold:
  1. i.
    The conjugate function is bounded:
    $\Big{|}G^{*}_{\tau}\big{(}Q(s,\cdot)\big{)}-\max_{a\in\mathcal{A}}Q(s,a)\Big{|}$ $\leq\tau B$
  2. ii.
    The gradient of the convex conjugate can be caharacterized as follows:
    $\nabla G_{\tau}^{*}(Q(s,\cdot))$ $=\arg\max_{\pi}\big{\{}\mathbb{E}_{a\sim\pi(\cdot|s)}[Q(s,a)]-G_{\tau}(\pi(\cdot|s))\big{\}}$
  3. iii.
    (Lipschitz gradients) For some constant $L_{G}>0$.
    $\left\|\nabla G_{\tau}^{*}(Q_{1}(s,\cdot))-\nabla G_{\tau}^{*}(Q_{1}(s,\cdot))\right\|$ $\leq\frac{L_{G}}{\tau}\left\|Q_{1}(s,\cdot)-Q_{2}(s,\cdot)\right\|$
  4. iv.
    The optimal policy $\pi^{*}_{\tau}(\cdot|s)=\nabla G^{*}_{\tau}\left(Q^{*}_{\tau}\left(s,\cdot\right)\right)$, for all $s\in\mathcal{S}$.